3(x/2+14)=180
I have no idea how to begin this equation.

Answer by stanbon(75887)

3(x/2+14)=180
------
Divide both sides by 3 to get:
x/2 + 14 = 60
---
Subtract 14 from both sides to get:
x/2 = 46
---
Multiply both sides by 2 to get:
x = 92
=============
Cheers,
Stan H.
=============